### v1

# Facilitating Automated Online Consensus Building through Parallel Thinking

Abstract.
Consensus building is inherently challenging due to the diverse opinions held by stakeholders. Effective facilitation is crucial to support the consensus building process and enable efficient group decision making. However, the effectiveness of facilitation is often constrained by human factors such as limited experience and scalability. In this research, we propose a Parallel Thinking-based Facilitation Agent (PTFA) that facilitates online, text-based consensus building processes. The PTFA automatically collects textual posts and leverages large language models (LLMs) to perform all of the six distinct roles of the well-established Six Thinking Hats technique in parallel thinking. To illustrate the potential of PTFA, a pilot study was carried out and PTFA’s ability in idea generation, emotional probing, and deeper analysis of ideas was demonstrated. Furthermore, a comprehensive dataset that contains not only the conversational content among the participants but also between the participants and the agent is constructed for future study.
collective intelligence, consensus building, automated facilitation agent, parallel thinking, six hats
$c$s: Human-centered computing Empirical studies in collaborative and social computing
1. Introduction
Collective intelligence, the emergent capability of groups to solve problems and make decisions collaboratively, has become a cornerstone of effective decision making in diverse domains, from business strategy to public policy (Becker et al., 2017). By integrating the knowledge, creativity, and perspectives of multiple individuals, collective intelligence can produce outcomes that surpass those achievable by individuals working in isolation. However, harnessing this potential often requires skilled facilitation to guide discussions, manage conflicts, and ensure equitable participation. To address the need for effective group consensus building facilitation, artificial intelligence (AI)-based approaches leveraging machine learning and intelligent agents have been proposed to support tasks such as content generation, balancing participation, and accumulating facilitation knowledge (Ito et al., 2022; Hadfi et al., 2023; Gu et al., 2021). In particular, the advent of generative AI and large language models (LLMs) has demonstrated significant potential in enhancing facilitation tasks, including collaborative brainstorming (Nomura et al., 2024), generating and refining statements (Tessler et al., 2024) and mitigating conflicts (Ma et al., 2024a). Despite these advances, achieving systematic automated facilitation support remains challenging. It requires abilities such as sophisticated contextual understanding, maintaining the natural flow of discussions, and striking a delicate balance between structured guidance and flexibility to avoid rigidity. Moreover, the significant energy cost of training LLMs and the popularity of well-known LLMs like ChatGPT makes it desirable to use pre-trained LLMs as well as necessary to investigate the capabilities of existing LLMs for such a specific application.
To bridge the gap between specific facilitation task support and systematic automated facilitation, we propose a novel Parallel Thinking-based Facilitation Agent (PTFA) designed to support online group consensus building. This conversational agent actively participates in text-based online discussions, guiding groups through structured, context-aware interactions to foster mutual understanding and agreement. It integrates the principles of parallel thinking, a structured methodology for systematically exploring multiple perspectives, with the advanced capabilities of large language models (LLMs) (Göçmen and Coşkun, 2019; Wan et al., 2024). Specifically, PTFA is based on the Six Thinking Hats technique, where each hat represents a role of an effective facilitator, e.g., the blue hat represents the role of managing the discussion and the green hat represents fostering creativity. The roles and suitable discussion interactions are selected automatically and dynamically based on the contributions of the participants. By dynamically adapting to the flow of discussions, PTFA adjusts its facilitation role in response to the evolving context, generating targeted facilitation inputs to enrich the discussion.
However, new and up-coming applications in this domain, like PTFA, will likely rely on readily-available LLMs, which require fine-tuned instructions based on their characteristics. The availability and popularity of certain LLMs, like ChatGPT and Gemini, means that, at least in the short term, they are likely to be used for LLM-based applications, including tools that are intended to support discussion facilitation. Hence, we based our PTFA implementation on a ChatGPT system prompt customization that uses the ChatGPT API to generate facilitator responses. By utilizing this approach we are able to provide insights that this approach has the potential to improve AI-based automatic facilitation, but needs specific tuning to the timeliness, phasing of discussion and consistency to provide an effective facilitator and a practical tool.
To evaluate PTFA and the general-purpose LLM approach, we developed an online text-based discussion platform, serving as an interactive interface for participants to gather and conduct discussions. We conducted a pilot study involving 16 discussion groups and 48 participants in an online environment collecting a rich dataset on participant interaction with the PTFA which has allowed us to establish guidelines on pursuing AI-based consensus-building platform development in idea generation, emotional probing and deeper argument analysis as well as challenges such as automated phase management and dynamic response timing control. Furthermore, this dataset will be public available when the full paper gets accepted.
2. Related Work
2.1. Group Decision Support with Automated Facilitation
Group decision making benefits significantly from facilitation, which can occur during, before, or after group discussions (Bostrom et al., 1993; Beranek et al., 1993; Zaraté et al., 2013). Traditionally, human facilitators have been essential in guiding discussions, requiring expertise in both the subject matter and the language of communication (Gimpel et al., 2024). However, the shortage of skilled facilitators creates a gap that modern automated facilitation systems can help address (Gu et al., 2021). Most existing automated facilitation approaches rely on rule-based and static reasoning systems (Gu et al., 2021). While these systems provide structure, their lack of adaptability limits their effectiveness in dynamic group interactions. Recently, LLMs have gained prominence as tools for brainstorming and consensus building, offering a dynamic and adaptive alternative. LLMs engage in active collaboration with human users, overcoming the rigidity of traditional systems (Noack et al., 2025; Wan et al., 2024; Joosten et al., 2024; Bouschery et al., 2023). LLMs have demonstrated promise in several aspects of group decision making. They facilitate agreement building within groups (Tessler et al., 2024; Shin et al., 2022), enhance the quality of discourse (Argyle et al., 2023), and contribute to better decision-making outcomes (Ma et al., 2024a). Additionally, LLMs can foster diversity of opinions (Kim et al., 2020) and improve inclusivity, such as by increasing women’s participation in discussions (Hadfi et al., 2023).
Despite these advances, the assessment of LLMs as generalized facilitators, capable of performing multiple tasks simultaneously while actively participating in conversations, remains underexplored. This paper aims to evaluate the effectiveness of LLMs in such a role, specifically through the lens of the Six Thinking Hats methodology (De Bono, 2017). Furthermore, we assess the quality of deliberations from the user’s perspective, a dimension that has been identified as lacking sufficient research (Shortall et al., 2022).
2.2. Facilitation with the Six Thinking Hats Method
The Six Thinking Hats framework is a well-established method for enhancing group decision-making and problem-solving. It involves six metaphorical hats, each representing a distinct mode of thinking. By focusing on one type of thinking at a time, the method reduces confusion and improves efficiency in group discussions (De Bono, 2017). The Six Thinking Hats methodology has been employed in experimental settings with encouraging results, particularly for creative brainstorming (Göçmen and Coşkun, 2019; Fender and Stickney, 2017). Previous studies have explored the use of this framework in facilitating human reflection on LLM-generated outputs, where participants evaluate and critique ideas rather than having LLMs actively assume the roles of the six hats (Zha et al., 2024).
In contrast, our approach assigns each LLM agent the role of one of the Six Thinking Hats to facilitate real-time discussions with human participants. While prior work has explored similar roles for LLMs in agent-to-agent discussions (Lu et al., 2024; Mushtaq et al., 2025), to the best of our knowledge this study is the first to evaluate their effectiveness in facilitating human group interactions through this method. The Six Thinking Hats approach can lead to higher-quality argumentation and foster critical thinking (Ekahitanond, 2018). However, it may also increase the cognitive load for participants, requiring careful consideration of its implementation (Ma et al., 2024b). By leveraging LLMs to assume these roles, we aim to mitigate the cognitive burden on human participants while enhancing the overall quality and depth of discussions. Our study is the first to assess the impact of using LLMs to actively embody the Six Thinking Hats roles in real-time group discussions, providing new insights into their potential for automated facilitation and collaborative decision-making.
3. Methodology
Overall: The discussion platform is developed using the Discourse forum system  [^1] https://www.discourse.org/, integrated with LLMs to facilitate structured conversations based on the Six Thinking Hats methodology (De Bono, 2017). The platform assists users in reaching consensus through structured interactions. The system consists of several key components, including a Discourse-based web interface where participants engage in discussions, a back-end database for storing discussion logs and participant interactions, and the OpenAI  [^2] https://openai.com/Assistants API, which provides real-time responses from the six thinking hats agents. WebSockets are used to enable real-time communication between the front-end and back-end components, while third-party plugins enhance forum functionalities such as logging and analytics.
In terms of technical implementation, the system utilizes the Discourse API to retrieve users’ post content and associated metadata (including timestamps, user IDs, and topic IDs), and the OpenAI Assistants API to generate responses based on the Six Thinking Hats framework. Discussion logs are securely stored in a PostgreSQL database, and participants are assigned unique identifiers to maintain anonymity. Security measures include HTTPS encryption to secure communications and role-based access control to manage permissions.
[…]
system
Figure 1. The Framework of PTFA.
Parallel Thinking-based Facilitation Agent Implementation: The facilitation agents in the system are implemented using OpenAI’s ChatGPT assistant API, which enables real-time responses to guide structured discussions based on the Six Thinking Hats framework (shown in Fig.1). Each thinking hat is represented by a distinct LLM agent that provides contextually appropriate responses. The White Hat agent provides factual data without interpretation, the Red Hat agent encourages the sharing of feelings and gut reactions, and the Black Hat agent identifies potential risks and downsides. The Yellow Hat agent highlights positive aspects and potential benefits, the Green Hat agent encourages innovative ideas and alternatives, and the Blue Hat agent directs the conversation and ensures structure.
The implementation of LLM-based facilitation agents requires careful prompt engineering at both a macro and micro level. At a macro level, prompts should be designed to provide comprehensive context by outlining the discussion background, defining the agent’s role, and specifying the overall objectives. This includes instructing the agent to guide participants towards consensus, foster balanced discussions, and adhere to structured communication principles. At the micro level, prompts should incorporate specific instructions that allow the agent to dynamically assess the current state of the conversation and intervene appropriately. These prompts should include mechanisms to determine when intervention is necessary, such as identifying signs of topic drift, unresolved conflicts, or opportunities for further elaboration. The prompts also define response styles, ensuring that the agent’s interventions are concise, relevant, and aligned with the discussion phase.
To minimize unnecessary interruptions and maintain a smooth discussion flow, a special mechanism is implemented in which the LLM outputs ’Good’ when it determines that the current conversation is progressing well without the need for further intervention. These outputs are automatically filtered and hidden from the discussion interface to avoid disrupting the participants. This approach ensures that the system provides support only when necessary, enhancing the overall user experience by reducing distractions while maintaining the integrity of the discussion.
Prompt engineering for these LLM-based facilitation agents is designed to ensure that each agent’s responses align with its designated role while maintaining clarity and conciseness. The primary prompts provide general guidance based on the respective thinking mode, whereas situational prompts encourage specific interventions tailored to the evolving discussion context. For instance, the White Hat prompts focus on factual accuracy and verification, while the Green Hat encourages brainstorming new ideas. The agents are triggered based on predefined timing intervals. Timing intervals are set based on the progression of the discussion phase, with interventions occurring at 30-second intervals to ensure continuous engagement without overwhelming the participants. If the system detects a lack of diversity in input or prolonged inactivity, the agents are prompted to re-engage users with targeted suggestions or requests for further elaboration.
4. User Study
4.1. Participants & Study Setting
To evaluate the efficacy of the proposed PTFA in supporting online consensus building, we held text-based real-world online discussions. Before embarking on this study, we secured ethical approval from the Faculty Ethics Committee at XX University (hidden for anonymization), guaranteeing adherence to established guidelines for ethical conduct in research involving human subjects. A total of 48 participants were recruited through lecture announcements with faculty permission, email invitations via department mailing lists, and posters placed in university common areas. 16 discussion groups were formed and each discussion group consisted of 3 participants, who were randomly selected while ensuring a balance in gender and English proficiency levels in each group. Detailed participant characteristics are shown in Table 1 in the appendix. Specifically, each group included at least two different genders and participants with at least two different levels of English proficiency. The experiments were conducted in university tutorial rooms equipped with laptops that connected to the online discussion platform we developed. Figure 3 and Figure 4 in the appendix show examples of experiment scenarios and the screenshot of the online discussion platform, respectively. Each discussion group participated in two online discussions, each featuring a different combination of discussion topics and facilitation models, detailed setting is shown in Table 1 in the appendix, and followed by a short facilitator evaluation survey.
Each discussion session lasted for 20 minutes. Two topics related to daily life were selected as the discussion subjects. Topic0 was ”Please decide one activity that you would like to do together” and Topic1 was ”Please decide one film that you would like to watch together”. Additionally, two facilitation models were designed for comparison. Facilitation Model 0 represented a traditional facilitation model, in which three facilitation messages were generated based on the discussion’s progress. Specific facilitation messages can be found in the appendix. In contrast, Facilitation Model 1 was developed using the proposed PTFA approach, introducing an alternative facilitation method for comparison.
4.2. Results
Our dataset consists of 32 discussions, 2 by each of the 16 groups, with each group consisting of 3 human participants. In total these discussions consist of 16,656 words in 1,669 posts, of which the facilitator contributed 3,459 words in 217 posts. Each post has a corresponding timestamp and, for the LLM-generated posts, which hat was used to generate the text. Analysis of the discussion transcripts reveals that while there are promising indications of the efficacy of PTFA-based consensus building, there are also challenges in utilizing an LLM to make timely, appropriate contributions that consistently advance the discussion in a productive manner. The post-discussion surveys support this observation. The survey responses indicate that participants had more mixed feelings about the performance of the PTFA facilitator on a number of discussion attributes, including the extent to which the facilitator helped consensus decision making (see Figure 2(b)), when compared to a baseline facilitation approach. Below, we first present the survey results that provide an overview of participant experiences. We then examine the types of constructive conversation pattern observed in the discussions before analyzing the problematic timing and formulation of comments that likely contributed to the lower satisfaction scores.
4.2.1. Survey Responses
The complementing surveys (see also Section 4.1) were designed to evaluate the level of support and capabilities the PTFA approach already provides, which, while highlighting a general positive view, emphasize how much problematic LLM patterns (see Section 4.2.3) can dampen the facilitation experience and thus the overall experience and outcomes.
Generally, the user experience with either facilitation model has been positive in terms of user experience (see Figure 2(a)) and reached consensus (see Figure 2(c)). However, despite positive contributions by the PTFA-based facilitator (Facilitation Model 1) towards being an effective and practical consensus building facilitator (see Section 4.2.2) negative effects (see Section 4.2.3) lead to mixed facilitator ratings. We observe this by the response to the questions ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ to which the participation could indicate their agreement with 7 the responses: ‘Strongly Agree’, ‘Agree’, ‘Somewhat Agree’, ‘Neutral’, ‘Somewhat Disagree’, ‘Disagree’ and ‘Strongly Disagree’. While participants react positively to Facilitation Model 0, as Figure 2(b) shows, the PTFA-based facilitator is rated much more diversely in the responses while still being positive to neutral overall.
The mixed performance of the PTFA facilitator also affects the user experience and their agreement with the consensus. As Figure 2(a) highlights, while the user experience is still positive, it is slightly dampened by some of the problematic LLM patterns (see Section 4.2.3). The user experience assessment is based on the question ‘How would you rate the user experience of the platform?’ to which the participation could indicate their satisfaction with 7 the responses: ‘Very Satisfied’, ‘Satisfied’, ‘Somewhat Satisfied’, ‘Neutral’, ‘Somewhat Unsatisfied’, ‘Unsatisfied’ and ‘Very Unsatisfied’. Moreover, the PTFA facilitator also dampens the agreement with the consensus, especially due to disrupting the consensus forming with the LLM’s difficulty to keep to specified phases (see Problematic Facilitation Patterns in Section 4.2.3). This consensus agreement assessment is based on the question ‘Do you agree with the consensus reached in this discussion?’ to which the participants could indicate their agreement with the same 7 responses as above.
(a) Survey responses for the question ‘How would you rate the user experience of the platform?’ rated into 7 possible responses as indicated in the legend in Figure 2(d).
(b) Survey responses for the question ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ rated into 7 possible responses as indicated in the legend in Figure 2(e).
(c) Survey responses for the question ‘Do you agree with the consensus reached in this discussion?’ rated into 7 possible responses as indicated in the legend in Figure 2(e).
(d) The legend for Figure 2(a)
(e) The legend for Figure 2(b) and Figure 2(c)
Figure 2. Survey responses for the questions around the rating of the facilitator models and the agreement with the resulting consensus. The legend in Figure 2(e) shows the possible responses used for the questions in Figure 2(b) and Figure 2(c). The 50% line divides the bars at 50% of participants.
[…]
Bar graph showing the share of survey responses for the question ‘How would you rate the user experience of the platform?’ rated into the 7 possible responses ‘Very Satisfied’, ‘Satisfied’, ‘Somewhat Satisfied’, ‘Neutral’, ‘Somewhat Unsatisfied’, ‘Unsatisfied’ and ‘Very Unsatisfied’ as well as bar graphs showing the share of survey responses for the questions ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ and ‘Do you agree with the consensus reached in this discussion?’ rated into the 7 possible responses ‘Strongly Agree’, ‘Agree’, ‘Somewhat Agree’, ‘Neutral’, ‘Somewhat Disagree’, ‘Disagree’ and ‘Strongly Disagree’.
4.2.2. Constructive Facilitation Patterns
We observed several instances in which the PTFA facilitator made productive contributions to the discussion that are distinct from the types of contributions provided by the human participants. We identify four types of useful intervention, each of which was contributed by a particular thinking hat. In the following, we describe these useful interventions. Example discussion excerpts of positive contributions are presented in the appendix.
Highlighting and exploring emotional responses of participants (Red Hat). The human participants mostly gave shallow opinions about suggestions; they typically said little more than whether or not they like an idea. By asking focused questions about the participants’ emotional reactions and feelings, the PTFA facilitator helped them make better informed decisions about their existing ideas and generate new ones better aligned with their preferences.
Identifying positive/negative qualities of suggestions that meet the participants needs (Yellow/Black Hat) Identifying strengths and weaknesses of suggestions, the PTFA facilitator provided a deeper analysis of proposals than that typically provided by the human participants independently. These comments highlighted particular features of ideas that aligned or conflicted with participants’ expressed interests. Contributions of this type usually expanded on a positive or negative comment made by a participant.
Providing factual information about the discussion topic (White Hat) The LLM is able to generate relevant facts learned from its training data. It did this during discussions either when participants expressed that they did not know some factual information (for example, the runtime of a film) or explicitly asked the facilitator for such information. While it is possible for LLMs to hallucinate details, almost all the facts it generated were correct. One exception was a comment that a theater production would open in the future, even though it had actually already started. Inaccuracies like this are due to the cut-off date of the data used to train the model rather than hallucination of facts.
Introducing new suggestions to the discussion (Green Hat) The PTFA facilitator regularly provided or extended suggestions. These suggestions were often well received by participants and then selected as the consensus decision. The LLM has a tendency to combine or extend prior suggestions - particularly when generating ideas for a shared activity. While these extensions were also sometimes well received, on other occasions the participants joked about how many components the facilitator was adding to the activities.
4.2.3. Problematic Facilitation Patterns
While the PTFA facilitator is capable of making useful contributions, these are offset by the many comments that were unhelpful, or even counterproductive. We observed two main categories of issues: poor management of the discussion phase and inconsistent response timing. Example discussion excerpts of problematic contributions are presented in the appendix.
Poor Phase Management
The most significant of these issues is in scheduling the phases of the consensus building discussion. The LLM is prompted to facilitate a divergent first discussion stage encouraging idea generation, followed by a convergent second stage of consensus formation. In practice, the facilitator stayed in the first stage throughout the discussions, with the green and blue hat personas consistently asking for further suggestions until the end of the 20 minute period, including when participants were trying to form a consensus or explicitly stated that they had reached a consensus they were happy with. It also asked the participants to generate new suggestions while they were already in the process of sharing new suggestions. Often, these inappropriate requests for more ideas were ignored by the participants. However, in one discussion a consensus was reached until the facilitator intervened with more suggestions. The new ideas split the participants’ preferences without enough time for them to discuss and agree on a new unanimous decision.
Hence, to build an effective facilitator based on LLMs, it is thus necessary to allow the facilitator to be aware of phases and times and fine-tune the facilitator to judge the state of the discussion from the participants’ perspective.
Inconsistent response timing
Another issue is that the PTFA facilitator sometimes did not respond for long periods of time. In one instance, it makes its first contribution 8 minutes into the session, with a comment “starting” the discussion - despite the participants already having generated many ideas. Similarly, it sometimes ignores questions explicitly asked of it. On the other hand, the PTFA facilitator may intervene too much - responding after every individual comment without allowing the participants to develop their ideas. These repeated messages are typically those asking for more suggestions, as described above.
Hence, it is necessary to be aware that LLMs can interact differently in seemingly similar situations and fine-tune the LLM’s intervention rate based on the state of the discussion, with respect of the content, and its previous interventions.
4.3. Discussion
This study provides insight in to both the potential and limitations of LLM-augmented discussions. Analysis of the 32 facilitated discussions shows that in a PTFA-based framework, LLMs can make valuable contributions that are otherwise neglected by human participants — particularly through idea generation, emotional probing, and deeper analysis of positive and negative qualities of ideas. These findings demonstrate the potential utility of LLMs in discussion facilitation, offering examples of how a PTFA could automate skilled facilitation to improve outcomes of consensus building. Despite this, the survey results reveal that users had more mixed opinions of the PTFA facilitator when compared to a well-received baseline facilitator on several discussion attributes. This is supported by the qualitative analysis of discussion transcripts, in which we identified many instances of the PTFA facilitator making unhelpful or inappropriate contributions to the discussion.
The timing of LLM interventions proved particularly problematic. Often, these contributions are inappropriate because they come at the wrong time in the discussion. For example, when it asks for suggestions at the end of the discussion or gives an emotional analysis of a user’s non-emotional comment. On the other hand, the PTFA facilitator is also sometimes absent from the discussion and misses moments when an intervention could prove useful, such as when asked a direct question by a participant or when the participants need to converge on a consensus decision. This highlights the need to investigate how to control LLM contributions in multi-participant conversations, an area less studied than back-and-forth interactions with a single other participant.
5. Conclusions
This paper explored the potential of AI agents in supporting consensus building processes in a natural conversation scenario. We introduced the Parallel Thinking-based Facilitation Agent (PTFA) that utilized parallel thinking-based strategies aligned with general-purpose LLMs. A real-world pilot user study, including 48 participants, was conducted, summarizing the PTFA’s ability along with the challenging issues. We not only highlight insights how LLMs are suitable but also noted some limitations and open challenges, thus providing directions for utilizing LLMs effectively for facilitation agents. Moreover, based on the user study, we constructed a novel dataset rich in conversational content among participants and between participants and PTFA. Future work will focus on addressing the open challenges identified in the user study.
Acknowledgements.

### v2

@@
-Facilitating Automated Online Consensus Building through Parallel Thinking
-Abstract.
+
+Abstract
 Consensus building is inherently challenging due to the diverse opinions held by stakeholders.
@@
 In this research, we propose a Parallel Thinking-based Facilitation Agent (PTFA) that facilitates online, text-based consensus building processes.
-The PTFA automatically collects textual posts and leverages large language models (LLMs) to perform all of the six distinct roles of the well-established Six Thinking Hats technique in parallel thinking.
-To illustrate the potential of PTFA, a pilot study was carried out and PTFA’s ability in idea generation, emotional probing, and deeper analysis of ideas was demonstrated.
+The PTFA automatically collects real-time textual input and leverages large language models (LLMs) to perform all six distinct roles of the well-established Six Thinking Hats technique in parallel thinking.
+To illustrate the potential of the agent, a pilot study was conducted, demonstrating its capabilities in idea generation, emotional probing, and deeper analysis of idea quality.
+Additionally, future open research challenges such as optimizing scheduling and managing behaviors in divergent phase are identified.
 Furthermore, a comprehensive dataset that contains not only the conversational content among the participants but also between the participants and the agent is constructed for future study.
-collective intelligence, consensus building, automated facilitation agent, parallel thinking, six hats
-$c$s: Human-centered computing Empirical studies in collaborative and social computing
-1. Introduction
-Collective intelligence, the emergent capability of groups to solve problems and make decisions collaboratively, has become a cornerstone of effective decision making in diverse domains, from business strategy to public policy (Becker et al., 2017).
+1 Introduction
+Collective intelligence, the emergent capability of groups to solve problems and make decisions collaboratively, has become a cornerstone of effective decision making in diverse domains, from business strategy to public policy [2].
 By integrating the knowledge, creativity, and perspectives of multiple individuals, collective intelligence can produce outcomes that surpass those achievable by individuals working in isolation.
-However, harnessing this potential often requires skilled facilitation to guide discussions, manage conflicts, and ensure equitable participation.
-To address the need for effective group consensus building facilitation, artificial intelligence (AI)-based approaches leveraging machine learning and intelligent agents have been proposed to support tasks such as content generation, balancing participation, and accumulating facilitation knowledge (Ito et al., 2022; Hadfi et al., 2023; Gu et al., 2021).
-In particular, the advent of generative AI and large language models (LLMs) has demonstrated significant potential in enhancing facilitation tasks, including collaborative brainstorming (Nomura et al., 2024), generating and refining statements (Tessler et al., 2024) and mitigating conflicts (Ma et al., 2024a).
-Despite these advances, achieving systematic automated facilitation support remains challenging.
-It requires abilities such as sophisticated contextual understanding, maintaining the natural flow of discussions, and striking a delicate balance between structured guidance and flexibility to avoid rigidity.
-Moreover, the significant energy cost of training LLMs and the popularity of well-known LLMs like ChatGPT makes it desirable to use pre-trained LLMs as well as necessary to investigate the capabilities of existing LLMs for such a specific application.
+However, harnessing this potential often requires skilled facilitation aimed at supporting decision making, such as guiding the process, promoting constructive interaction and inclusivity, and managing content in a structured manner [9].
+To address the need for effective group consensus building facilitation, artificial intelligence (AI)-based approaches leveraging machine learning and intelligent agents have been proposed to support tasks such as content generation, balancing participation, and accumulating facilitation knowledge [13, 12, 10].
+In particular, the advent of generative AI and large language models (LLMs) has demonstrated significant potential in enhancing specific facilitation tasks, including collaborative brainstorming [22], generating and refining statements [25] and mitigating conflicts [17].
+Despite recent advances, achieving automated systematic facilitation support remains challenging, particularly because it involves managing multiple facilitation tasks concurrently and managing the relationships between those tasks in dynamic decision-making processes.
+For example, systematic facilitation requires the ability to understand the context of the decision-making process, determine how to promote participation, and adapt accordingly based on the identified characteristics and needs of the participants.
+Additionally, it must determine the appropriate timing for executing these facilitation tasks.
+This requires abilities such as sophisticated contextual understanding, maintaining the natural flow of discussions, and striking a delicate balance between structured guidance and flexibility to avoid rigidity.
 To bridge the gap between specific facilitation task support and systematic automated facilitation, we propose a novel Parallel Thinking-based Facilitation Agent (PTFA) designed to support online group consensus building.
-This conversational agent actively participates in text-based online discussions, guiding groups through structured, context-aware interactions to foster mutual understanding and agreement.
-It integrates the principles of parallel thinking, a structured methodology for systematically exploring multiple perspectives, with the advanced capabilities of large language models (LLMs) (Göçmen and Coşkun, 2019; Wan et al., 2024).
-Specifically, PTFA is based on the Six Thinking Hats technique, where each hat represents a role of an effective facilitator, e.g., the blue hat represents the role of managing the discussion and the green hat represents fostering creativity.
+This conversational agent actively participates in the group decision-making process, guiding stakeholders through structured, context-aware interactions to foster mutual understanding and consensus.
+To address multiple facilitation tasks concurrently, PTFA integrates the principles of parallel thinking, a structured methodology for systematically exploring multiple perspectives, with the advanced capabilities of LLMs [11, 26].
+Specifically, PTFA is based on the Six Thinking Hats technique, which is a successful method for finding consensus [6].
+Each hat represents a role of an effective facilitator, e.g., the blue hat represents the role of managing the discussion and the green hat represents fostering creativity.
 The roles and suitable discussion interactions are selected automatically and dynamically based on the contributions of the participants.
-By dynamically adapting to the flow of discussions, PTFA adjusts its facilitation role in response to the evolving context, generating targeted facilitation inputs to enrich the discussion.
-However, new and up-coming applications in this domain, like PTFA, will likely rely on readily-available LLMs, which require fine-tuned instructions based on their characteristics.
-The availability and popularity of certain LLMs, like ChatGPT and Gemini, means that, at least in the short term, they are likely to be used for LLM-based applications, including tools that are intended to support discussion facilitation.
-Hence, we based our PTFA implementation on a ChatGPT system prompt customization that uses the ChatGPT API to generate facilitator responses.
-By utilizing this approach we are able to provide insights that this approach has the potential to improve AI-based automatic facilitation, but needs specific tuning to the timeliness, phasing of discussion and consistency to provide an effective facilitator and a practical tool.
-To evaluate PTFA and the general-purpose LLM approach, we developed an online text-based discussion platform, serving as an interactive interface for participants to gather and conduct discussions.
+To evaluate PTFA, we developed an online text-based discussion platform, serving as an interactive interface for participants to gather and conduct discussions.
 We conducted a pilot study involving 16 discussion groups and 48 participants in an online environment collecting a rich dataset on participant interaction with the PTFA which has allowed us to establish guidelines on pursuing AI-based consensus-building platform development in idea generation, emotional probing and deeper argument analysis as well as challenges such as automated phase management and dynamic response timing control.
-Furthermore, this dataset will be public available when the full paper gets accepted.
-2. Related Work
-2.1. Group Decision Support with Automated Facilitation
-Group decision making benefits significantly from facilitation, which can occur during, before, or after group discussions (Bostrom et al., 1993; Beranek et al., 1993; Zaraté et al., 2013).
-Traditionally, human facilitators have been essential in guiding discussions, requiring expertise in both the subject matter and the language of communication (Gimpel et al., 2024).
-However, the shortage of skilled facilitators creates a gap that modern automated facilitation systems can help address (Gu et al., 2021).
-Most existing automated facilitation approaches rely on rule-based and static reasoning systems (Gu et al., 2021).
+To summarize, the main contributions of this paper are as follows:
+  • We propose a novel agent-based framework that implements a multi-role facilitation model powered by an LLM, enabling automated and systematic support for group decision-making.
+  • We implement the proposed approach in an online, free-text decision-making scenario and conduct real-world pilot experiments to evaluate its potential in idea generation, emotional probing, and deeper analysis of idea quality and find limitations in optimizing scheduling and managing behaviors as a fully automated facilitator.
+  • We construct a novel dataset rich in conversational interactions, capturing dialogues both among participants and between participants and the agent.
+The remainder of this paper is organized as follows.
+Section 2 reviews related work on automated facilitation, conversational agents, and parallel thinking methodologies.
+Section 3 introduces the overall framework of the proposed agent-based facilitation system.
+Section 4 details the implementation and experimental setup, and presents the results of the pilot experiments, along with a discussion of their implications, limitations, and potential directions for future work.
+Finally, Section 5 concludes the paper.
+2 Related Work
+2.1 Group Decision Support with Automated Facilitation
+Group decision making benefits significantly from facilitation, which can occur during, before, or after group discussions [4, 3, 27].
+Traditionally, human facilitators have been essential in guiding discussions, requiring expertise in both the subject matter and the language of communication [9].
+However, the shortage of skilled facilitators creates a gap that modern automated facilitation systems are increasingly positioned to address [10].
+Most existing automated facilitation approaches rely on rule-based and static reasoning systems [10].
 While these systems provide structure, their lack of adaptability limits their effectiveness in dynamic group interactions.
 Recently, LLMs have gained prominence as tools for brainstorming and consensus building, offering a dynamic and adaptive alternative.
-LLMs engage in active collaboration with human users, overcoming the rigidity of traditional systems (Noack et al., 2025; Wan et al., 2024; Joosten et al., 2024; Bouschery et al., 2023).
-LLMs have demonstrated promise in several aspects of group decision making.
-They facilitate agreement building within groups (Tessler et al., 2024; Shin et al., 2022), enhance the quality of discourse (Argyle et al., 2023), and contribute to better decision-making outcomes (Ma et al., 2024a).
-Additionally, LLMs can foster diversity of opinions (Kim et al., 2020) and improve inclusivity, such as by increasing women’s participation in discussions (Hadfi et al., 2023).
-Despite these advances, the assessment of LLMs as generalized facilitators, capable of performing multiple tasks simultaneously while actively participating in conversations, remains underexplored.
-This paper aims to evaluate the effectiveness of LLMs in such a role, specifically through the lens of the Six Thinking Hats methodology (De Bono, 2017).
-Furthermore, we assess the quality of deliberations from the user’s perspective, a dimension that has been identified as lacking sufficient research (Shortall et al., 2022).
-2.2. Facilitation with the Six Thinking Hats Method
+In contrast to rigid systems, LLMs engage in active collaboration with human users, supporting more natural facilitation [21, 26, 14, 5].
+LLMs have demonstrated promise in several aspects of group decision-making.
+They facilitate agreement building within groups [25, 23], enhance the quality of discourse [1], and contribute to better decision-making outcomes [17].
+Additionally, LLMs can foster diversity of opinions [15] and improve inclusivity, such as by increasing women’s participation in discussions [12].
+However, prior LLM-based facilitation approaches tend to focus on technical optimization, or surface level coordination [19], overlooking research in other fields that offer structured methods for enhancing group discussion.
+We argue that effective automated facilitation should be grounded in established methodologies that are designed to support reflective, balanced, and goal-oriented discourse.
+To this end, this paper explores the role of LLMS as generalized facilitators, able to perform different tasks and roles, through the lens of the Six Thinking Hats methodology [6].
+This structured approach to parallel thinking is well-established in the organizational literature but has not been utilized in automated facilitation.
+Furthermore, we assess the quality of deliberations from the user’s perspective, a dimension that has been identified as lacking sufficient research [24].
+2.2 Facilitation with the Six Thinking Hats Method
 The Six Thinking Hats framework is a well-established method for enhancing group decision-making and problem-solving.
 It involves six metaphorical hats, each representing a distinct mode of thinking.
-By focusing on one type of thinking at a time, the method reduces confusion and improves efficiency in group discussions (De Bono, 2017).
-The Six Thinking Hats methodology has been employed in experimental settings with encouraging results, particularly for creative brainstorming (Göçmen and Coşkun, 2019; Fender and Stickney, 2017).
-Previous studies have explored the use of this framework in facilitating human reflection on LLM-generated outputs, where participants evaluate and critique ideas rather than having LLMs actively assume the roles of the six hats (Zha et al., 2024).
+By focusing on one type of thinking at a time, the method reduces confusion and improves efficiency in group discussions [6].
+The Six Thinking Hats methodology has been employed in experimental settings with encouraging results, particularly for creative brainstorming [11, 8].
+Previous studies have explored the use of this framework in facilitating human reflection on LLM-generated outputs, where participants evaluate and critique ideas rather than having LLMs actively assume the roles of the six hats [28].
 In contrast, our approach assigns each LLM agent the role of one of the Six Thinking Hats to facilitate real-time discussions with human participants.
-While prior work has explored similar roles for LLMs in agent-to-agent discussions (Lu et al., 2024; Mushtaq et al., 2025), to the best of our knowledge this study is the first to evaluate their effectiveness in facilitating human group interactions through this method.
-The Six Thinking Hats approach can lead to higher-quality argumentation and foster critical thinking (Ekahitanond, 2018).
-However, it may also increase the cognitive load for participants, requiring careful consideration of its implementation (Ma et al., 2024b).
+While prior work has explored similar roles for LLMs in agent-to-agent discussions [16, 20], to the best of our knowledge, this study is the first to evaluate their effectiveness in facilitating human group interactions through this method.
+The Six Thinking Hats approach can lead to higher-quality argumentation and foster critical thinking [7].
+However, it may also increase the cognitive load for participants, requiring careful consideration of its implementation [18].
 By leveraging LLMs to assume these roles, we aim to mitigate the cognitive burden on human participants while enhancing the overall quality and depth of discussions.
 Our study is the first to assess the impact of using LLMs to actively embody the Six Thinking Hats roles in real-time group discussions, providing new insights into their potential for automated facilitation and collaborative decision-making.
-3. Methodology
-Overall: The discussion platform is developed using the Discourse forum system  [^1] https://www.discourse.org/, integrated with LLMs to facilitate structured conversations based on the Six Thinking Hats methodology (De Bono, 2017).
-The platform assists users in reaching consensus through structured interactions.
-The system consists of several key components, including a Discourse-based web interface where participants engage in discussions, a back-end database for storing discussion logs and participant interactions, and the OpenAI  [^2] https://openai.com/Assistants API, which provides real-time responses from the six thinking hats agents.
-WebSockets are used to enable real-time communication between the front-end and back-end components, while third-party plugins enhance forum functionalities such as logging and analytics.
-In terms of technical implementation, the system utilizes the Discourse API to retrieve users’ post content and associated metadata (including timestamps, user IDs, and topic IDs), and the OpenAI Assistants API to generate responses based on the Six Thinking Hats framework.
-Discussion logs are securely stored in a PostgreSQL database, and participants are assigned unique identifiers to maintain anonymity.
-Security measures include HTTPS encryption to secure communications and role-based access control to manage permissions.
-[…]
-system
-Figure 1. The Framework of PTFA.
-Parallel Thinking-based Facilitation Agent Implementation: The facilitation agents in the system are implemented using OpenAI’s ChatGPT assistant API, which enables real-time responses to guide structured discussions based on the Six Thinking Hats framework (shown in Fig.1).
+3 Methodology
+Overall Implementation.
+To support the PTFA framework, we developed an online discussion platform based on the Discourse forum system [^1] https://www.discourse.org/, enhanced with large language model (LLM) integration.
+The platform enables structured, real-time text-based discussions, with LLM agents acting as facilitators according to the Six Thinking Hats methodology.
+The system architecture consists of a web-based interface for user participation, a backend database to store user-generated content and metadata, and an integration with the OpenAI Assistants API [^2] https://openai.com/ for dynamic response generation.
+Communication between frontend and backend components is handled using WebSockets to ensure real-time responsiveness.
+To manage discussions securely and effectively, each participant is assigned a unique anonymous identifier, and all discussion data is stored in a PostgreSQL database.
+The platform employs HTTPS encryption for secure communication and role-based access control to restrict system access.
+In addition, third-party plugins are used to support logging, moderation, and analytics, providing additional functionality for monitoring and evaluating discussion quality.
+The LLM agents are triggered based on predefined time intervals or inactivity detection, ensuring responsive but non-intrusive facilitation throughout the conversation.
+Figure 1: The framework of PTFA.
+Parallel Thinking-based Facilitation Agent Implementation.
+The facilitation agents in the system are implemented using OpenAI’s ChatGPT assistant API, which enables real-time responses to guide structured discussions based on the Six Thinking Hats framework (shown in Fig.1).
 Each thinking hat is represented by a distinct LLM agent that provides contextually appropriate responses.
-The White Hat agent provides factual data without interpretation, the Red Hat agent encourages the sharing of feelings and gut reactions, and the Black Hat agent identifies potential risks and downsides.
-The Yellow Hat agent highlights positive aspects and potential benefits, the Green Hat agent encourages innovative ideas and alternatives, and the Blue Hat agent directs the conversation and ensures structure.
 The implementation of LLM-based facilitation agents requires careful prompt engineering at both a macro and micro level.
@@
 The prompts also define response styles, ensuring that the agent’s interventions are concise, relevant, and aligned with the discussion phase.
+The prompt engineering for each of the six hat agents is detailed below:
+  • White Hat (Facts and Information): This agent embodies objective thinking.
+Its prompt instructs it to focus solely on data and facts without interpretation, grounding the conversation in reality.
+It is designed to intervene when the discussion requires objective information or when participants make unsupported statements.
+An example of a situational prompt is: "Could you clarify the exact figures or facts related to this issue?
+Here’s what we know so far: [insert relevant data]."
+  • Red Hat (Emotions and Intuition): Representing feelings and gut reactions, this agent allows for the expression of emotion without justification.
+The primary prompt directs it to provide a short, intuitive perspective when the conversation touches on personal feelings.
+A situational prompt for this agent is: "This feels like an emotional moment.
+How are we feeling about this issue right now?"
+  • Black Hat (Critical Thinking and Risk): This agent provides a critical, cautious perspective.
+Its core instruction is to identify potential risks, weaknesses, and flaws that the group might be overlooking, ensuring negative outcomes are considered.
+An example of its situational prompt is: "Have we considered the potential downsides?
+Here’s a risk we might be overlooking: [insert risk]."
+  • Yellow Hat (Optimism and Benefits): The Yellow Hat agent focuses on optimism and positive thinking.
+It is prompted to offer a constructive view by highlighting the benefits and potential advantages of ideas, especially when the discussion needs a boost of optimism.
+A corresponding situational prompt is: "Looking at the bright side, this idea offers some exciting opportunities we shouldn’t overlook."
+  • Green Hat (Creativity and Alternatives): This agent encourages creative and innovative thinking.
+It is prompted to intervene with new ideas or alternative approaches when the conversation becomes stalled or could benefit from novel solutions.
+An example of a creative situational prompt is: "What if we approached this from a different angle?
+Here’s an idea to consider: [insert new idea]."
+  • Blue Hat (Process and Control): The Blue Hat agent manages the thinking process itself.
+Its instructions are to provide structure and direction, organize the conversation, and ensure the discussion stays on track and that all perspectives are considered.
+A situational prompt for managing the process is: "It seems like we’re getting off track.
+Maybe we should focus on this key point: [insert key point]."
 To minimize unnecessary interruptions and maintain a smooth discussion flow, a special mechanism is implemented in which the LLM outputs ’Good’ when it determines that the current conversation is progressing well without the need for further intervention.
@@
 The agents are triggered based on predefined timing intervals.
-Timing intervals are set based on the progression of the discussion phase, with interventions occurring at 30-second intervals to ensure continuous engagement without overwhelming the participants.
+Timing intervals are set based on the progression of the discussion phase, with interventions to ensure continuous engagement without overwhelming the participants.
 If the system detects a lack of diversity in input or prolonged inactivity, the agents are prompted to re-engage users with targeted suggestions or requests for further elaboration.
-4. User Study
-4.1. Participants & Study Setting
+4 User Study
+4.1 Participants & Study Setting
 To evaluate the efficacy of the proposed PTFA in supporting online consensus building, we held text-based real-world online discussions.
-Before embarking on this study, we secured ethical approval from the Faculty Ethics Committee at XX University (hidden for anonymization), guaranteeing adherence to established guidelines for ethical conduct in research involving human subjects.
+Before embarking on this study, we secured ethical approval from the Faculty Ethics Committee at University of Southampton, guaranteeing adherence to established guidelines for ethical conduct in research involving human subjects.
 A total of 48 participants were recruited through lecture announcements with faculty permission, email invitations via department mailing lists, and posters placed in university common areas.
 16 discussion groups were formed and each discussion group consisted of 3 participants, who were randomly selected while ensuring a balance in gender and English proficiency levels in each group.
-Detailed participant characteristics are shown in Table 1 in the appendix.
 Specifically, each group included at least two different genders and participants with at least two different levels of English proficiency.
 The experiments were conducted in university tutorial rooms equipped with laptops that connected to the online discussion platform we developed.
-Figure 3 and Figure 4 in the appendix show examples of experiment scenarios and the screenshot of the online discussion platform, respectively.
-Each discussion group participated in two online discussions, each featuring a different combination of discussion topics and facilitation models, detailed setting is shown in Table 1 in the appendix, and followed by a short facilitator evaluation survey.
+Figure 2 shows the screenshot of the online discussion platform.
+Each discussion group participated in two online discussions, each featuring a different combination of discussion topics and facilitation models and followed by a short facilitator evaluation survey.
+The purpose of the survey was to evaluate the participants’ perception of the facilitator’s effectiveness, overall user experience and the degree to which consensus was achieved.
+By comparing the responses between the facilitator models, we aim to evaluate the practical impact of PTFA on group discussion dynamics and outcomes.
+Through this process we identify areas of improvement for LLM-assisted facilitation.
+Figure 2: Screenshot of online discussion platform
 Each discussion session lasted for 20 minutes.
-Two topics related to daily life were selected as the discussion subjects.
-Topic0 was ”Please decide one activity that you would like to do together” and Topic1 was ”Please decide one film that you would like to watch together”.
+To alleviate barriers to participation, two topics related to daily life were selected as discussion subjects.
+Topic0 was "Please decide one activity that you would like to do together" and Topic1 was "Please decide one film that you would like to watch together".
 Additionally, two facilitation models were designed for comparison.
-Facilitation Model 0 represented a traditional facilitation model, in which three facilitation messages were generated based on the discussion’s progress.
-Specific facilitation messages can be found in the appendix.
+Facilitation Model 0 represented a traditional facilitation model in which three facilitation messages were generated, each after a certain amount of elapsed time:
+  • “Hi all, our goal today is to reach a consensus on the question posed at the end of the discussion. Please start by generating ideas.” was posted at the beginning of the discussion,
+  • “You have already discussed it for 10 mins. This is a good time for you to reconsider the ideas that you have already had.” was posted 10 minutes after the discussion began, and
+  • “There are only 3 minutes left, if you haven’t reached a consensus yet, please make a decision as soon as possible.” was posted 17 minutes after the discussion started.
 In contrast, Facilitation Model 1 was developed using the proposed PTFA approach, introducing an alternative facilitation method for comparison.
-4.2. Results
-Our dataset consists of 32 discussions, 2 by each of the 16 groups, with each group consisting of 3 human participants.
-In total these discussions consist of 16,656 words in 1,669 posts, of which the facilitator contributed 3,459 words in 217 posts.
+Intervention decisions are made at 30-second intervals, taking group size into account.
+4.2 Results
+Our study consisted of 48 participants, with 3 in each of the 16 groups.
+Each group participated in two discussion giving a dataset of 32 discussions containing 16,656 words in 1,669 posts, of which the facilitator contributed 3,459 words in 217 posts.
 Each post has a corresponding timestamp and, for the LLM-generated posts, which hat was used to generate the text.
+Participants were almost entirely aged between 18 and 34, with a near even split between males and females.
+Most participants (68.8%) considered themselves a fluent or native speaker of english, with only 2.1% considering themselves a beginner, the least proficient category.
+A full breakdown of the participant demographics is presented in Figure 3.
+(a) The share of the participants by their ages.
+(b) The share of the participants by their genders.
+(c) The share of the participants by their proficiencies in English.
+Figure 3: The characteristics of the participants in terms of age, gender and English proficiency.
 Analysis of the discussion transcripts reveals that while there are promising indications of the efficacy of PTFA-based consensus building, there are also challenges in utilizing an LLM to make timely, appropriate contributions that consistently advance the discussion in a productive manner.
@@
 Below, we first present the survey results that provide an overview of participant experiences.
-We then examine the types of constructive conversation pattern observed in the discussions before analyzing the problematic timing and formulation of comments that likely contributed to the lower satisfaction scores.
-4.2.1. Survey Responses
+We then examine the types of constructive conversation patterns observed in the discussions before analyzing the problematic timing and formulation of comments that likely contributed to the lower satisfaction scores.
+4.2.1 Survey Responses
 The complementing surveys (see also Section 4.1) were designed to evaluate the level of support and capabilities the PTFA approach already provides, which, while highlighting a general positive view, emphasize how much problematic LLM patterns (see Section 4.2.3) can dampen the facilitation experience and thus the overall experience and outcomes.
 Generally, the user experience with either facilitation model has been positive in terms of user experience (see Figure 4(a)) and reached consensus (see Figure 4(c)).
-However, despite positive contributions by the PTFA-based facilitator (Facilitation Model 1) towards being an effective and practical consensus building facilitator (see Section 4.2.2) negative effects (see Section 4.2.3) lead to mixed facilitator ratings.
-We observe this by the response to the questions ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ to which the participation could indicate their agreement with 7 the responses: ‘Strongly Agree’, ‘Agree’, ‘Somewhat Agree’, ‘Neutral’, ‘Somewhat Disagree’, ‘Disagree’ and ‘Strongly Disagree’.
-While participants react positively to Facilitation Model 0, as Figure 2(b) shows, the PTFA-based facilitator is rated much more diversely in the responses while still being positive to neutral overall.
+However, despite positive contributions by the PTFA-based facilitator (Facilitation Model 1) towards being an effective and practical consensus building facilitator (see Section 4.2.2), negative effects (see Section 4.2.3) lead to mixed facilitator ratings.
+We observe this in the response to the question ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ to which the participant could indicate their agreement with one of the statements: ‘Strongly Agree’, ‘Agree’, ‘Somewhat Agree’, ‘Neutral’, ‘Somewhat Disagree’, ‘Disagree’ and ‘Strongly Disagree’.
+While participants reacted more positively to the PTFA-based facilitator - being slightly above 50% - in comparison to Facilitation Model 0 - being slightly below 50% - as Figure 4(b) shows, overall, the rating of both is mixed.
 The mixed performance of the PTFA facilitator also affects the user experience and their agreement with the consensus.
@@
 The user experience assessment is based on the question ‘How would you rate the user experience of the platform?’ to which the participation could indicate their satisfaction with 7 the responses: ‘Very Satisfied’, ‘Satisfied’, ‘Somewhat Satisfied’, ‘Neutral’, ‘Somewhat Unsatisfied’, ‘Unsatisfied’ and ‘Very Unsatisfied’.
-Moreover, the PTFA facilitator also dampens the agreement with the consensus, especially due to disrupting the consensus forming with the LLM’s difficulty to keep to specified phases (see Problematic Facilitation Patterns in Section 4.2.3).
+Moreover, the PTFA facilitator also slightly dampens the agreement with the consensus, especially due to disrupting the consensus forming with the LLM’s difficulty to keep to specified phases (see 7 in Section 4.2.3).
 This consensus agreement assessment is based on the question ‘Do you agree with the consensus reached in this discussion?’ to which the participants could indicate their agreement with the same 7 responses as above.
@@
 (e) The legend for Figure 4(b) and Figure 4(c)
-Figure 2.
-Survey responses for the questions around the rating of the facilitator models and the agreement with the resulting consensus.
+Figure 4: Survey responses for the questions around the rating of the facilitator models and the agreement with the resulting consensus.
 The legend in Figure 4(e) shows the possible responses used for the questions in Figure 4(b) and Figure 4(c).
 The 50% line divides the bars at 50% of participants.
-[…]
-Bar graph showing the share of survey responses for the question ‘How would you rate the user experience of the platform?’ rated into the 7 possible responses ‘Very Satisfied’, ‘Satisfied’, ‘Somewhat Satisfied’, ‘Neutral’, ‘Somewhat Unsatisfied’, ‘Unsatisfied’ and ‘Very Unsatisfied’ as well as bar graphs showing the share of survey responses for the questions ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ and ‘Do you agree with the consensus reached in this discussion?’ rated into the 7 possible responses ‘Strongly Agree’, ‘Agree’, ‘Somewhat Agree’, ‘Neutral’, ‘Somewhat Disagree’, ‘Disagree’ and ‘Strongly Disagree’.
-4.2.2. Constructive Facilitation Patterns
-We observed several instances in which the PTFA facilitator made productive contributions to the discussion that are distinct from the types of contributions provided by the human participants.
+4.2.2 Constructive Facilitation Patterns
+We observed cases in which the PTFA facilitator made productive contributions to the discussion that were distinct from the types of contributions provided by the human participants.
 We identify four types of useful intervention, each of which was contributed by a particular thinking hat.
 In the following, we describe these useful interventions.
-Example discussion excerpts of positive contributions are presented in the appendix.
+Example discussion excerpts of positive contributions are presented in Figure 5 and Figure 6.
+(a) The red hat externalizing its analysis of the participants comment.
+(b) The white hat responding with accurate factual information to a question posed by a participant.
+Figure 5: Examples of constructive facilitator contributions
+(a) The yellow hat highlighting positives of a movie. The white hat providing accurate factual information about a movie in response to a participant expressing their unfamiliarity with it.
+(b) The yellow hat highlighting positive aspects of ideas that resonate with the participants.
+Figure 6: Examples of constructive facilitator contributions
 Highlighting and exploring emotional responses of participants (Red Hat).
-The human participants mostly gave shallow opinions about suggestions; they typically said little more than whether or not they like an idea.
+The human participants mostly gave shallow opinions about suggestions; they typically said little more than whether or not they liked an idea.
 By asking focused questions about the participants’ emotional reactions and feelings, the PTFA facilitator helped them make better informed decisions about their existing ideas and generate new ones better aligned with their preferences.
-Identifying positive/negative qualities of suggestions that meet the participants needs (Yellow/Black Hat) Identifying strengths and weaknesses of suggestions, the PTFA facilitator provided a deeper analysis of proposals than that typically provided by the human participants independently.
+Identifying positive/negative qualities of suggestions that meet the participants needs (Yellow/Black Hat).
+Identifying strengths and weaknesses of suggestions, the PTFA facilitator provided a deeper analysis of proposals than that typically provided by the human participants independently.
 These comments highlighted particular features of ideas that aligned or conflicted with participants’ expressed interests.
 Contributions of this type usually expanded on a positive or negative comment made by a participant.
-Providing factual information about the discussion topic (White Hat) The LLM is able to generate relevant facts learned from its training data.
-It did this during discussions either when participants expressed that they did not know some factual information (for example, the runtime of a film) or explicitly asked the facilitator for such information.
+Providing factual information about the discussion topic (White Hat).
+The LLM is able to generate relevant facts learned from its training data.
+It did this during discussions, either when participants expressed that they did not know some factual information (for example, the runtime of a film) or explicitly asked the facilitator for such information.
 While it is possible for LLMs to hallucinate details, almost all the facts it generated were correct.
 One exception was a comment that a theater production would open in the future, even though it had actually already started.
-Inaccuracies like this are due to the cut-off date of the data used to train the model rather than hallucination of facts.
-Introducing new suggestions to the discussion (Green Hat) The PTFA facilitator regularly provided or extended suggestions.
-These suggestions were often well received by participants and then selected as the consensus decision.
-The LLM has a tendency to combine or extend prior suggestions - particularly when generating ideas for a shared activity.
-While these extensions were also sometimes well received, on other occasions the participants joked about how many components the facilitator was adding to the activities.
-4.2.3. Problematic Facilitation Patterns
-While the PTFA facilitator is capable of making useful contributions, these are offset by the many comments that were unhelpful, or even counterproductive.
+Inaccuracies like this are due to the cut-off date of the data used to train the model rather than the hallucination of facts.
+Introducing new suggestions to the discussion (Green Hat).
+The PTFA facilitator regularly provided or extended suggestions.
+These suggestions were often well-received by participants and then selected as the consensus decision.
+The LLM has a tendency to combine or extend prior suggestions, particularly when generating ideas for a shared activity.
+While these extensions were also sometimes well received, on other occasions, the participants joked about how many components the facilitator was adding to the activities.
+4.2.3 Problematic Facilitation Patterns
+While the PTFA facilitator is capable of making useful contributions, these are offset by the many comments that were unhelpful or even counterproductive.
 We observed two main categories of issues: poor management of the discussion phase and inconsistent response timing.
-Example discussion excerpts of problematic contributions are presented in the appendix.
-Poor Phase Management
+Example discussion excerpts of problematic contributions are presented in Figure 7.
+(a) Asking for idea suggestions repeatedly and at an inappropriate time. These comments begin 4 minutes before the end of the 20 minute discussion period.
+(b) The red hat externalizing its analysis of a participant’s comment.
+Figure 7: Examples of problematic facilitator behaviors.
+Poor Phase Management.
 The most significant of these issues is in scheduling the phases of the consensus building discussion.
-The LLM is prompted to facilitate a divergent first discussion stage encouraging idea generation, followed by a convergent second stage of consensus formation.
-In practice, the facilitator stayed in the first stage throughout the discussions, with the green and blue hat personas consistently asking for further suggestions until the end of the 20 minute period, including when participants were trying to form a consensus or explicitly stated that they had reached a consensus they were happy with.
+The LLM is prompted to facilitate a divergent first discussion stage, encouraging idea generation, followed by a convergent second stage of consensus formation.
+In practice, the facilitator stayed in the first stage throughout the discussions, with the green and blue hat personas consistently asking for further suggestions until the end of the 20-minute period, including when participants were trying to form a consensus or explicitly stated that they had reached a consensus they were happy with.
 It also asked the participants to generate new suggestions while they were already in the process of sharing new suggestions.
 Often, these inappropriate requests for more ideas were ignored by the participants.
-However, in one discussion a consensus was reached until the facilitator intervened with more suggestions.
+However, in one discussion, a consensus was reached until the facilitator intervened with more suggestions.
 The new ideas split the participants’ preferences without enough time for them to discuss and agree on a new unanimous decision.
-Hence, to build an effective facilitator based on LLMs, it is thus necessary to allow the facilitator to be aware of phases and times and fine-tune the facilitator to judge the state of the discussion from the participants’ perspective.
-Inconsistent response timing
+Hence, to build an effective facilitator based on LLMs, it is necessary to allow the facilitator to be aware of phases and times and fine-tune the facilitator to judge the state of the discussion from the participants’ perspective.
+Inconsistent response timing.
 Another issue is that the PTFA facilitator sometimes did not respond for long periods of time.
@@
 Similarly, it sometimes ignores questions explicitly asked of it.
-On the other hand, the PTFA facilitator may intervene too much - responding after every individual comment without allowing the participants to develop their ideas.
+On the other hand, the PTFA facilitator may intervene too much, responding after every individual comment without allowing the participants to develop their ideas.
 These repeated messages are typically those asking for more suggestions, as described above.
-Hence, it is necessary to be aware that LLMs can interact differently in seemingly similar situations and fine-tune the LLM’s intervention rate based on the state of the discussion, with respect of the content, and its previous interventions.
-4.3. Discussion
+Hence, it is necessary to be aware that LLMs can interact differently in seemingly similar situations and fine-tune the LLM’s intervention rate based on the state of the discussion, with respect to the content, and its previous interventions.
+4.3 Discussion
 This study provides insight into both the potential and limitations of LLM-augmented discussions.
-Analysis of the 32 facilitated discussions shows that in a PTFA-based framework, LLMs can make valuable contributions that are otherwise neglected by human participants — particularly through idea generation, emotional probing, and deeper analysis of positive and negative qualities of ideas.
+Analysis of the 32 facilitated discussions shows that in a PTFA-based framework, LLMs can make valuable contributions that are otherwise neglected by human participants, particularly through idea generation, emotional probing, and deeper analysis of positive and negative qualities of ideas.
 These findings demonstrate the potential utility of LLMs in discussion facilitation, offering examples of how a PTFA could automate skilled facilitation to improve outcomes of consensus building.
@@
 This highlights the need to investigate how to control LLM contributions in multi-participant conversations, an area less studied than back-and-forth interactions with a single other participant.
-5. Conclusions
+5 Conclusions
 This paper explored the potential of AI agents in supporting consensus building processes in a natural conversation scenario.
 We introduced the Parallel Thinking-based Facilitation Agent (PTFA) that utilized parallel thinking-based strategies aligned with general-purpose LLMs.
-A real-world pilot user study, including 48 participants, was conducted, summarizing the PTFA’s ability along with the challenging issues.
-We not only highlight insights how LLMs are suitable but also noted some limitations and open challenges, thus providing directions for utilizing LLMs effectively for facilitation agents.
+A real-world pilot user study, including 48 participants, was conducted, demonstrating the PTFA’s ability while also identifying a number of remaining challenges.
+We not only highlight insights on how LLMs are suitable but also note some limitations and open challenges, thus providing directions for utilizing LLMs effectively for facilitation agents.
 Moreover, based on the user study, we constructed a novel dataset rich in conversational content among participants and between participants and PTFA.
 Future work will focus on addressing the open challenges identified in the user study.
-Acknowledgements.
+{credits}
+5.0.1 Acknowledgements
+This work was supported by JST CREST JPMJCR20D1, and JSPS KAKENHI Grant Number JP24K20900.
+Additional support for this work was provided by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (EP/V022067/1) on Citizen-Centric AI Systems (https://ccais.ac.uk/) and through the FEVER Programme Grant (EP/W005883/1).